Daily Events on 2025/04/17 (Thu)

last updated at  2025/04/17 16:45:09 JST


<< prev day

@

next day >>


Station

Pass No.
20250417-

Contact time (UT)

Max. Elev.
 (deg)

Events

AOS

LOS

MSP1

8000

08:07:44

08:20:59

50.7

OP/OG upload
XRT: TI-CMD
EIS: TI-CMD

USC20

0300

09:13:19

09:26:22

66.8

OP/OG upload, XRT:TI-CMD, EIS:TI-CMD (backup)
AOCS: Upload pointing table
EIS: Table Upload
DR playback (by OP)

SNT1

4010

10:04:00

10:16:32

26.2

XRT: Table Upload

USC20

0400

10:52:40

11:01:08

8.1

No-commanding pass
DR playback (by OP)

USC34

0500

21:13:23

21:25:35

21.9

DR playback (by OP)

USC34

0600

22:50:50

23:02:37

25.5

DR playback (by OP)

Flare list(LM) : http://www.lmsal.com/solarsoft/last_events/


DR information

OP main: 2025/04/17 11:32:00 - 2025/04/19 10:58:00 (UT)

Remaining amount
(prediction)

4299 Mbits @ OP_START ( 2025/04/17 11:32:00 )


DR usage: The DR would become empty on Apr 18 at 10 UT if all telescopes took no images until that time.


SOT

XRT

EIS

DR allocation [%]

54

23

23

DR allocation [Mbits]

3356

1429

1429

DR TLMLMT [Mbits]

1302 @OP_START
-> 5576

2350

2350


daily note

EIS:
- Support HOP 412
- AR 14060 obs

XRT:
- HOP 412
- HOP 349 & 448
- AR 14060 obs

SOT:
- HOP 412
- HOP 393
- AR 14060 obs

POINTING:
2025/04/17 12:20:30.000  |    -21  |   -985  |  HOP412  01/20
2025/04/17 12:43:30.000  |    -21  |   -885  |  HOP412  02/20
2025/04/17 13:06:30.000  |    -21  |   -785  |  HOP412  03/20
2025/04/17 13:59:00.000  |    -21  |   -685  |  HOP412  04/20
2025/04/17 14:22:00.000  |    -21  |   -585  |  HOP412  05/20
2025/04/17 14:45:00.000  |    -21  |   -485  |  HOP412  06/20
2025/04/17 15:39:00.000  |    -21  |   -385  |  HOP412  07/20
2025/04/17 16:02:00.000  |    -21  |   -285  |  HOP412  08/20
2025/04/17 16:25:00.000  |    -21  |   -185  |  HOP412  09/20
2025/04/17 17:27:00.000  |    -21  |    -85  |  HOP412  10/20
2025/04/17 17:50:00.000  |    -21  |      5  |  HOP412  11/20
2025/04/17 19:03:30.000  |    -21  |    105  |  HOP412  12/20
2025/04/17 19:26:30.000  |    -21  |    205  |  HOP412  13/20
2025/04/17 20:40:00.000  |    -21  |    305  |  HOP412  14/20
2025/04/17 21:03:00.000  |    -21  |    405  |  HOP412  15/20
2025/04/17 21:26:00.000  |    -21  |    505  |  HOP412  16/20
2025/04/17 22:17:00.000  |    -21  |    605  |  HOP412  17/20
2025/04/17 22:40:00.000  |    -21  |    705  |  HOP412  18/20
2025/04/17 23:03:00.000  |    -21  |    805  |  HOP412  19/20
2025/04/17 23:52:00.000  |    -21  |    905  |  HOP412  20/20
2025/04/18 00:15:00.000  |      0  |      0  |  End of HOP412
EOP

[Maintenance Note (Yamagishi), for CO/CP/RT Tohban]
Some software packages will be updated and security patch will be installed on machines
'aocswin' and 'higashirinkan' and 'ope-gw'
at 15:00-16:30(JST) on 17th April. When that's over, these machines will be rebooted.
Please do not use these machines during updating, and report if you have any trouble.


CHIEF
TOHBAN

SUB
TOHBAN

CP

SOT CO

XRT CO

EIS CO

EVENING

Kotanagi

Toriumi

Toriumi

Hirose

Brooks/McKevitt

MORNING


operation files

pointing file : cmdpln/20250417/re-point_202504178000.txt